OpenStudy (anonymous):

the bottom of a valley is 10 m below sea level. a nearby city is 120 m above sea level. what is the difference in elevations between the valley floor and the city

jimthompson5910 (jim_thompson5910):

"10 m below sea level." is the number -10 "120 m above sea level." is the number +120

jimthompson5910 (jim_thompson5910):

subtract the two +120 - (-10) = 120 + 10 = 130 so the distance between the two cities, the vertical distance, is 130 meters